### v1

# QTIP: Quantization with Trellises and Incoherence Processing

Abstract
Post-training quantization (PTQ) reduces the memory footprint of LLMs by quantizing weights to low-precision datatypes. Since LLM inference is usually memory-bound, PTQ methods can improve inference throughput. Recent state-of-the-art PTQ approaches have converged on using vector quantization (VQ) to quantize multiple weights at once, which improves information utilization through better shaping. However, VQ requires a codebook with size exponential in the dimension. This limits current VQ-based PTQ works to low VQ dimensions ($\leq 8$) that in turn limit quantization quality. Here, we introduce QTIP, which instead uses trellis coded quantization (TCQ) to achieve ultra-high-dimensional quantization. TCQ uses a stateful decoder that separates the codebook size from the bitrate and effective dimension. QTIP introduces a spectrum of lookup-only to computed lookup-free trellis codes designed for a hardware-efficient “bitshift” trellis structure; these codes achieve state-of-the-art results in both quantization quality and inference speed.
1 Introduction
Large language models (LLMs) have accelerated advancements in fields ranging from natural language processing [34] to scientific modeling [28]. However, the largest LLMs have hundreds of billions of parameters that can take over a terabyte of memory to load in half-precision; this size poses significant challenges for the practical deployment of LLMs [33, 18, 2]. For example, small-batch autoregressive decoding, a common form of inference for LLMs, is memory bound [35]. Even on a modern datacenter GPU with $\approx 3$TB/s memory bandwidth, a large LLM ($\gtrapprox 200$GB) can only be directly run at $\lessapprox 20$ tokens per second and requires multiple devices [4]. One way to accelerate inference is by compressing LLMs. This directly reduces the memory footprint of the model and increases the theoretical maximum inference throughput on any given machine.
One form of compression, weight-only post-training quantization (PTQ), quantizes trained model weights to lower precision datatypes [10, 35, 6]. The latest state-of-the-art weight-only PTQ methods, QuIP##\## and AQLM, use vector quantization (VQ) to achieve high-quality 2-bit models [35, 12]. In VQ, a vector $x\in\mathbb{R}^{d}$ is quantized to one of $2^{kd}$ vectors in $\mathbb{R}^{d}$ that form a codebook $C\in\mathbb{R}^{2^{kd}\times d}$. A higher vector dimension $d$ allows for better codebook shaping and packing density, improving information utilization [20]. However, VQ requires exponential time and space in both the bitrate and dimension, limiting its practicality. During quantization, VQ costs $O(2^{kd}d)$ time to perform nearest-neighbor rounding to $C$, and during inference, $C$ must fit in hardware cache for fast lookups. This exponential scaling limits how high $d$ can be and thus the advantages of VQ over scalar quantization.
To address this limitation, we propose QTIP, which uses trellis-coded quantization (TCQ) to enable tractable ultra-high-dimensional ($>100$) quantization and improves quantization quality over prior VQ-based approaches. In the simplest scalar form of TCQ, a length-$T$ sequence $S$ is statefully quantized using a trellis – a directed graph $G$ with $2^{L}$ nodes, each with $2^{k}$ incoming and outgoing edges and a scalar value [24]. The reconstructed sequence $\hat{S}$ corresponds to the node values of a length-$T$ walk on $G$, and quantization finds the walk that minimizes some distortion metric on $S$ and $\hat{S}$. Since neighboring entries in $\hat{S}$ are connected by one of $2^{k}$ edges, we only need to store which edge an entry came from, which takes $k$ bits. For additive distortion metrics such as squared error, the optimal $\hat{S}$ can be found with the Viterbi algorithm, which runs in $O(2^{L}T)$ time [14, 24]. This means that the cost of quantization is independent of the bitrate $k$ and linear in the sequence dimension $T$, enabling tractable high dimensional quantization.
However, TCQ is not free. During inference, vanilla TCQ requires storing both $G$ and the size $2^{L}\times V$ node value codebook, which can be too large to fit in cache. TCQ-quantized sequences also cannot generally be decoded in parallel, as $t$th elment of $\hat{S}$ could depend on up to the first $tk$ encoded bits. In QTIP, we solve these issues by introducing a series of fast compute-based Gaussian codes designed for the hardware-efficient “bitshift trellis.” Specifically, the bitshift trellis supports parallel decoding, does not require storing $G$, and our compute-based codes eliminate needing to store a large node value codebook. This enables high-quality quantization of Gaussian sources while supporting fast inference, and we adopt incoherence processing with the random Hadamard transform to ensure that LLM weights are approximately i.i.d Gaussian distributed. Altogether, QTIP
  1. 1.
    Simultaneously achieves state-of-the-art weight-only LLM PTQ quality and fast inference through a combination of hardware-efficient trellis and codebook design.
  2. 2.
    Introduces multiple novel hardware-efficient ($\leq 4$ instructions per weight) computed and hybrid lookup-computed random Gaussian codes for TCQ on i.i.d. Gaussian sources.
Refer to caption     Refer to caption
Figure 1: QTIP performs ultra-high dimensional ($>100$) quantization by using Trellis Coded Quantization, which has linear cost in dimension. This enables QTIP to outperform Vector Quantization-based approaches (QuIP##\##, AQLM) that are limited to small dimensions.
2 Background and Related Works
We focus on weight-only post-training quantization (PTQ) of LLMs in this work; other model-compression approaches include quantization-aware training (QAT) and pruning. These methods are not strictly orthogonal to each other, as one could both prune and quantize a model. Since QTIP is a weight-only PTQ method, the rest of this section focuses on this area. Most current state-of-the-art PTQ methods round to minimize the per-layer proxy loss from Nagel et al. [27].
$\ell(\hat{W})=\mathbb{E}_{x}\left[\|(\hat{W}-W)x\|^{2}\right]=\operatorname{tr}\left((\hat{W}-W)H(\hat{W}-W)^{T}\right)$ (1)
Here, $\hat{W}\in\mathbb{R}^{m\times n}$ is the quantized weight matrix, $x\in\mathbb{R}^{n}$ is an input activation, and $H=\mathbb{E}_{x}\left[xx^{T}\right]\in\mathbb{R}^{n\times n}$ is interpreted as a proxy Hessian matrix. This objective is defined per-layer, making it tractable for very large models. However, minimizing it is difficult due to the non-differentiable nature of quantization. Instead many works have proposed algorithms such as Hessian-based adaptive rounding, alternating optimization, and even coordinate descent to approximately minimize the proxy error [12, 6, 35, 15].
2.1 Incoherence Processing
The effectiveness of these methods depends on properties of $W$. For example, many works have observed that weight and activation outliers cause poor quantization quality [11, 21, 30]. In QuIP, Chee et al. [6] proposed that incoherence was important for quantifying this effect.
Definition 2.1 (Chee et al. [6]).
A Hessian $H\in\mathbb{R}^{n\times n}$ is $\mu$-incoherent if its eigendecomposition $H=Q\Lambda Q^{T}$ has $\max_{i,j}\;|Q_{ij}|=\max_{i,j}\;|e_{i}^{T}Qe_{j}|\leq\mu/\sqrt{n}$. A weight matrix $W\in\mathbb{R}^{m\times n}$ is $\mu$-incoherent if $\max_{i,j}\;\textstyle|W_{ij}|=\max_{i,j}\;|e_{i}^{T}We_{j}|\leq\mu\|W\|_{F}/\sqrt{mn}$.
Essentially, incoherence means the weights and important rounding directions (Hessian eigenvectors) are not too large in any direction, aiding quantization. To make $W,H$ incoherent (small $\mu$), one can perform incoherence processing (IP) by conjugating $W,H$ with random orthogonal matrices $U,V:\tilde{W}\leftarrow UWV^{T},\tilde{H}\leftarrow VHV^{T}$. QuIP##\## introduced IP with the random Hadamard transformation (RHT), which performs $\tilde{W}\leftarrow V_{m}S_{m}WS_{n}V_{n}^{T},\tilde{H}\leftarrow V_{n}S_{n}HS_{n}V_{n}^{T}$ where $V_{k}$ is a $k\times k$ Hadamard matrix and $S_{k}$ is a length $k$ random sign vector. The RHT achieves, with probability $\geq 1-\delta$, $\mu_{\tilde{W}}=2\log(4mn/\delta)$, meaning that $\tilde{W}$’s entries are approximately independently Gaussian distributed, which can aid quantization [35, 3]. We choose to build on incoherence processing here because the independent Gaussian-like weights it produces are suitable inputs for trellis coding [23].
2.2 Vector Quantization (VQ) for LLM PTQ
$k$-bit VQ quantizes a $d$ dimensional vector $S$ to one of $2^{kd}$ $d$-dimensional vectors that form a codebook $C\in\mathbb{R}^{2^{kd}\times d}$ [1]. Since $C$ is an unstructured collection of arbitrary vectors, VQ enables better shaping and packing density than scalar product quantization (SPQ), where each entry in $S$ is quantized independently [20]. However, this also comes at the cost of exponential time quantization and exponential space inference: finding the nearest neighbor in $C$ requires $O(2^{kd}d)$ time, and storing $C$ requires $O(2^{kd}d)$ space. The current crop of state-of-the-art LLM PTQ methods, QuIP##\## and AQLM, both use VQ to achieve high-quality 2-bit models. Since the shaping advantage of VQ comes from high dimensionality, both QuIP##\## and AQLM attempt to maximize dimensionality. AQLM’s uses a large 8D codebook (1MiB) that does not fit in L1 cache. QuIP##\## uses an 8D compressible codebook based on the $E_{8}$ lattice, which is highly symmetric. This codebook is compressible by 256×256$\times$256 × and barely fits in L1 cache. In either case, the VQ dimension is effectively hardware-limited to $\leq 8$, motivating methods that enable even higher-dimensional quantization.
2.3 Trellis-Coded Quantization (TCQ)
TCQ was first proposed by Marcellin and Fischer [24] to apply the benefits of trellis coded modulation, a conceptually dual problem, to quantization. Define a $(L,k,V)$ trellis $G$ as a directed graph with $2^{L}$ nodes, each of which has $2^{kV}$ incoming and outgoing edges and a value $\in\mathbb{R}^{V}$; these values form a codebook $C\in\mathbb{R}^{2^{L}\times V}$. To quantize a length-$T$ sequence $S\in\mathbb{R}^{T}$, each contiguous length-$V$ subsequence of $S$ is assigned to a node $\in G$, with the restriction that the assigned nodes form a walk. The reconstruction $\hat{S}$ of $S$ is then given by concatenating node values in the walk. When $V=1$, this setup describes Marcellin and Fischer [24]’s original scalar TCQ. When $V>1$, this describes TCVQ, which applies TCQ to vectors [13, 37].
Finding the optimal $\hat{S}$ under an additive distortion metric can be done with the Viterbi algorithm in $O(2^{L}T)$ time. This is linear in sequence length, enabling ultra-high dimensional quantization. For exposition, we briefly describe the Viterbi algorithm here. Concretely, if we want to quantize a $T$-length scalar sequence reinterpreted as a sequence of vectors $s_{1},s_{2},\ldots,s_{T/V}\in\mathbb{R}^{V}$ using a trellis code with graph $G$ and codebook $C$, this corresponds to solving the optimization problem
$\mbox{minimize}\;\sum_{i=1}^{T/V}\|C_{x_{i}}-s_{i}\|^{2}\;\;\mbox{ over }x_{1},x_{2},\ldots,x_{T/V}\mbox{ the vertex sequence of a walk on graph }G.$
This optimization problem can be solved exactly with dynamic programming via the value function
$\mathcal{V}_{t}(x)=\min\;\left\{\sum_{i=1}^{t}\|C_{x_{i}}-s_{i}\|^{2}\;\middle|\;x_{1},x_{2},\ldots,x_{t}\mbox{ the vertex sequence of a walk on }G\mbox{ and }x_{t}=x\right\}$
using the update rule
$\mathcal{V}_{t}(y)=\min_{(x,y)\in G}\mathcal{V}_{t-1}(x)+\|C_{y}-s_{t}\|^{2}.$
This Viterbi approach clearly takes time linear in $T$ and in the number of edges of $G$; with a few simple optimizations this can be brought to $O(2^{L}T)$. In comparison, brute-force-searching all possible $2^{kT}$ codes—which is what we would need to do for an unstructured $k$-bit $T$-dimensional codebook— would take time proportional to $2^{LT/V}$. The ability to tractably find the closest representable vector in $\mathbb{R}^{T}$, even for large $T$, is in some sense the “main benefit” of trellis coding. For i.i.d sources, as $L$ increases, TCQ efficiently approaches the infinite-length distortion-rate $D_{R}$, which lower bounds the attainable distortion of a $k$-bit quantizer [20]. As shown in Table 1, when quantizing an i.i.d. Gaussian with $k=2$, the scalar Lloyd-Max quantizer attains 0.118 MSE, QuIP##\##’s 8D E8P codebook 0.089 MSE, our (QTIP) 256D $L=16$ TCQ quantizer 0.069 MSE, and $D_{R}=0.063$ [22, 25, 35, 9].
3 QTIP
Quantizing with TCQ requires storing both the codebook ($2^{L}\times V$) and trellis structure ($2^{L}\times 2^{k}V$) during inference. These components are too large for fast inference when $L\gtrapprox 12$, which is necessary for high quality. Furthermore, for a generic trellis, recovering the state (and so the decoded value) at step $t$th requires a graph walk using the first $kt$ bits: this prevents parallel decoding. QTIP solves these problems with a novel combination of incoherence processing, a hardware-efficient “bitshift trellis,” and fast compute-based random Gaussian codes. Incoherence processing makes $W$ approximatelly i.i.d Gaussian, which reduces quantization to Gaussian source coding. The bitshift trellis removes needing to store the trellis structure during decoding and also enables parallel decoding. Finally, the fast compute-based random Gaussian codes remove the need to store the full codebook, completing the equation for fast inference. On the quality side, the fast random Gaussian codes enable the simple bitshift trellis to match complicated trellises and achieve state-of-the-art quantization quality.
The main focus of QTIP is on what to quantize with (i.e. TCQ) and not how to quantize (e.g. adaptive rounding or descent methods). The general construction of QTIPcan be used as a drop-in replacement for VQ in any rounding framework. In the following sections, we first describe the “bitshift” trellis (Section 3.1). Then, we describe a series of fast compute-based codes for i.i.d Gaussian sources, aligning with different types of hardware (Sections 3.1.1 and 3.1.2). Finally, we give an approximation for the tail-biting trellis problem, which lets us more efficiently load weights in hardware (Section 3.2).
3.1 “Bitshift” Trellis and Codebook Design
Figure 2: A bitshift trellis code with $L=2,k=1,V=1$. Nodes 0, 1, 2, and 3 have code values 0.5, 0.1, 0.8, and 0.3, respectively. Each node can only transition to the $22^{kV}=$ 2 nodes that share their top $L-kV=1$ bit with its bottom $L-kV=1$ bit. In this example, $\hat{S}$ can be stored as 0010110. $\hat{S}$ is also tail-biting, so the last $L-kV=1$ bits can be dropped to give $\hat{S}=$ 001011.
The bitshift trellis was introduced by Mao and Gray [23] as part of the “random permutation trellis coder” (RPTC). In the bitshift trellis, node $i$ has an edge to node $j$ if $\exists c\in\mathbb{Z},0\leq c<2^{kV},$ s.t. $j=(i2^{kV}\mbox{ mod }2^{L})+c$. Essentially, the top $L-kV$ bits of $j$ equal the bottom $L-kV$ bits of $i$. This means that the first group of $V$ weights depends only on the bits at positions $\{1,2,\ldots,L\}$, the second only on bit positions $\{kV+1,kV+2,\ldots,kV+L\}$, and in general the $t$th on bit positions $\{(t-1)kV+1,\ldots,(t-1)kV+L\}$. During inference, obtaining the next compressed group of $V$ weights in a sequence only requires bitshifting by $kV$ bits, which is supported on virtually all hardware. Furthermore, since each group of $V$ weights only depends on a contiguous window of $L$ bits in $\hat{S}$, decoding can be parallelized. Figure 2 shows a simple $(L=2,k=1,V=1)$ bitshift trellis. Note that edges only exist between nodes that overlap by 1 bit, and storing the quantized length 6 $\hat{S}$ indeed only requires 6 bits (plus the initial state).
Quantizing an i.i.d. source with the bitshift trellis is nontrivial because neighboring groups of weights sharing many bits can potentially lead to strong correlations (Figure 3 LL). The RPTC permutes the codebook to decorrelate neighboring weight groups (Figure 3 RR) [23]. However, this requires storing the codebook or storing and applying the permutation, both of which are prohibitively expensive during decoding. Instead, QTIP introduces a series of compute-based codes to produce a psuedorandom code, which has the same decorrelating effect and admits fast inference. To match approximately i.i.d. Gaussian RHT-transformed matrices, these codes produce psuedorandom approximate Gaussians in as few as 2 instructions per weight (see Table 1 and Figure 3). To the best of our knowledge, these code constructions alone are novel and we are the first to propose a lookup-free Gaussian trellis code.
Table 1: QTIP’s compute-based codes (1MAD, 3INST, HYB) achieve similar distortion rates as a pure-lookup random Gaussian trellis code (RPTC) when quantizing an i.i.d Gaussian source to 2 bits. All TCQ methods ($L=16$) outperform SQ and VQ and are significantly closer to the infinite-length distortion rate $D_{R}$, which lower bounds the distortion a $k$-bit quantizer can attain.
SQ VQ 1D TCQ 2D TCQ
Quant. Lloyd-Max QuIP##\## E8P 1MAD 3INST RPTC HYB RPTC $D_{R}$
Dim. 1 8 256 256 256 256 256 $\infty$
MSE. 0.118 0.089 0.069 0.069 0.068 0.071 0.069 0.063
3.1.1 Lookup-Free Computed Codes
Figure 3: Set of representable neighboring values in a bitshift trellis with $L=16,k=2,V=1$ for (far left) a code with strong correlations, (left center) algorithm 1 (“1MAD”), (right center) algorithm 2 (“3INST”), and (far right) a random Gaussian code. Note that while 1MAD has minor correlations, both 1MAD and 3INST are close to a random Gaussian, resulting in good quantization quality.
Here, we present two pure-computed lookup-free codes that produce a pseudorandom approximately Gaussian number from a $L$ bit word, enabling fast decoding on cache-limited hardware. These codes avoid strong correlations and can be implemented in $\leq 4$ hardware instructions per weight on NVIDIA GPUs. We present two codes here to illustrate that multiple such codes are possible: in practice a lookup-free code can be designed to use the instructions available on whatever hardware we want to run on.
Algorithm 1 (1MAD) first runs a linear congruential generator (LCG) to produce a pseudorandom 32-bit word. This requires 2 instructions (MAD and &). It then sums the 32-bit word as four 8-bit unsigned integers; this sum is approximately Gaussian distributed. This requires 1 instruction (vabsdiff4). Finally, this sum must be scaled and shifted (another MAD). Although there are only $2^{10}$ representable values even when $L>10$, this does not empirically affect quantization quality. 1MAD requires choosing $a$ and $b$ to avoid strong correlations; we set $a=34038481$ and $b=76625530$ (Figure 3 LC).
Algorithm 2 (3INST) also first runs an LCG to produce a random 32-bit word $X$. Then, it XORs the bottom 16 bits of $X$ with the mantissa bits, bottom two exponent bits, and sign bit of a magic FP16 number $m$ to produce an FP16 number $m_{1}$. It then repeats this with the top 16 bits of $X$ to produce $m_{2}$ and returns $m_{1}+m_{2}$. This entire process can be implemented in 3 ALU instructions [^1] As there is currently no instruction on NVIDIA GPUs that sums the top and bottom half of a 32-bit word as two FP16s, this requires an extra data movement instruction to “split” the 32-bit word into two 16-bit registers. with a MAD for the LCG, a lop3 to mask and XOR with a packed duplicated $m$, and then summing $m_{1}$ and $m_{2}$. $m_{1}+m_{2}$ is approximately distributed by the sum of two mirrored exponential distributions, which is close to Gaussian. Like with Algorithm 1, $a,b,$ and $m$ must be chosen to avoid correlations; we used $a=89226354,b=64248484,m=0.922$ (Figure 3 right).
Algorithm 1 Computed Gaussian Code “1MAD”
0:  $L$-bit 0 left-padded integer $x$, uint32 $a,b$.
  $x\leftarrow(ax+b)\mbox{\ mod\ }2^{32}$run LCG to get uniform random $x$} {sum $x$ as four 8-bit unsigned integers, this is approximately Gaussian}
[…]
  $x\leftarrow(x-510)/147.8$
  Pseudorandom approximate Gaussian $x$.
Algorithm 2 Computed Gaussian Code “3INST”
0:  $L$-bit 0 left-padded integer $x$, uint32 $a,b$, float16 $m$.
  $x\leftarrow(ax+b)\mbox{\ mod\ }2^{32}$run LCG to get uniform random $x$} {modify sign, mantissa, and bottom 2 exponent bits of $m$ and sum, this is approximately Gaussian}
[…]
  Pseudorandom approximate Gaussian $x$.
3.1.2 Hybrid Lookup-Computed Codes
Here, we describe a hybrid computed-lookup code that computes a pseudorandom (or hashed) index into a 2D vector codebook ($V=2$). This code is tailored for modern GPUs, which have enough cache for a small in-memory LUT—one benefit of using such a LUT over a purely computed codebook is that a LUT can be fine-tuned after quantization. Algorithm 3 first performs the hash $X\leftarrow X^{2}+X$ to mix the lower order and upper order bits of $X$ [19]. Then, it takes bits $(14-Q+1)-14$ (0 indexed) as an index into a $2^{Q}\times 2$ LUT to get two 16-bit floats. (The reason why we chose a 2D codebook here is that shared memory on NVIDIA GPUs is accessed in 32-bit-word elements, and each such word can contain two 16-bit floats.) Finally, it XORs bit 15 of $X$ to flip the sign of the second entry of the codebook vector. Algorithm 3 can be implemented with MAD, bitshift, mask, and lop3, giving an amortized 2 instructions per weight. This effectively assigns a $L$ bit word to one of $2^{Q+1}$ 2D vectors, each of which can be fine-tuned to improve quality. Algorithm 3 can also be implemented to XOR bit 31 alongside bit 15 (this is free in the lop3) to give an effectively $2^{Q+2}$-sized codebook, which can improve quantization quality. We only realized this after running all the experiments, so the numbers in this paper use the “one sign flip” version of Algorithm 3. In QTIP, we initialize the LUT using K-means on an empirical 2D i.i.d. Gaussian distribution.
Algorithm 3 Hybrid Computed-Lookup 2D Gaussian Code “HYB”
0:  $L$-bit 0 left-padded integer $x$, codebook $C\in\mathbb{R}^{2^{Q}\times(V=2)}$.
  $x\leftarrow x\cdot x+x\mbox{\ mod\ }2^{32}$calculate hash}
[…]
  Pseudorandom approximate Gaussian vector $v$.
3.2 Tail-Biting Trellises
Directly quantizing a length-$T$ sequence to a $(L,k,V)$ trellis results in a total of $kT+L-kV$ bits since the starting state takes an additional $L-kV$ bits to store. If we run inference on a machine with $w$-bit words where $w|kT$, we must read an extra $\lceil\frac{L-kV}{w}\rceil w-(L-kV)$ wasted bits per sequence. For common $w$ (e.g. 32), setting $L=kV+w$ makes the Viterbi algorithm intractable. One way to solve this is by enforcing that the start and end state share $L-kV$ bits, i.e. the trellis is tail-biting [5]. Exactly solving the tail-biting trellis problem via dynamic programming takes time quadratic in the state space ($2^{L}$), making this problem intractable for reasonable $L\geq 12$ [31]. However, since RHT-processed weights are approximately i.i.d., simple algorithms can be effective for approximately solving the tail-biting problem. We propose Algorithm 4, which first rotates the sequence by $T/2$, quantizes it, and then extracts the overlap between the rotated start and end states. It then requantizes the original sequence with this overlap as the tail-biting overlap. This only requires two Viterbi calls in total. Table 2 shows that in practice, Algorithm 4 can find close-to-optimal tail-biting sequences while being significantly cheaper to run than other tail-biting approximation algorithms [31].
Algorithm 4 Tail-biting Trellis Approx. 0:  Sequence $S\in\mathbb{R}^{T}$, $(L,k,V)$ Trellis $G$.   $S^{\prime}\leftarrow$ Rotate $S$ to the right by $\lfloor T/2\rfloor$   $\hat{S}^{\prime}\leftarrow$ Viterbi($S^{\prime}$, $G$)   $O\leftarrow L-kV$ bit overlap of $\hat{S}^{\prime}_{\lfloor T/2\rfloor}\hat{S}^{\prime}_{\lfloor T/2\rfloor+1}$   $\hat{S}\leftarrow$ Viterbi($S$, $G$) with start/end overlap $=O$   Tail biting $\hat{S}$
Table 2: Quantizing 4K $T=256$ i.i.d Gaussian seqs. with a tail-biting $(12,k,1)$ trellis.
$k$ Alg. 4 MSE Optimal MSE
1 0.2803 0.2798
2 0.0733 0.0733
3 0.0198 0.0198
4 0.0055 0.0055
4 Experiments
Here, we present experiments quantizing the Llama family of models with QTIP [33, 34, 26]. These models offer strong performance across a wide range of sizes, allowing us to compare how different quantization methods perform and scale. We primarily compare QTIP against QuIP##\## and AQLM. For Llama 1, we include GPTVQ-2D instead of AQLM since AQLM does not publish Llama 1 numbers [36]. GPTVQ-2D performs 2D VQ inside GPTQ and offers strong performance. These methods outperform scalar quantization methods including GPTQ, AWQ, and OmniQuant; comparisons to those methods can be found in QuIP##\## and AQLM [21, 15, 30, 35, 12]. We mainly focus on the hybrid code (Section 4.2) since it is tailored for modern GPUs, and present a full suite of results for it. For the computed codes (Section 4.1), we present results for Llama 2.
Since the proxy error is not an additive distortion metric, we cannot minimize it by quantizing $W$ as one sequence. Instead, for all experiments, we use QTIP as a quantizer in QuIP##\##’s BlockLDLQ, which allows us to simultaneously achieve high dimensionality and low proxy error [35]. Specifically, we quantize a block of $T_{x}\times T_{y}$ weights as a sequence, where $T_{x}$ and $T_{y}$ span the output and input dimensions of $W$, respectively. Since BlockLDLQ only specifies feedback along the input dimension, this is equivalent to BlockLDLQ with $g=T_{y}$ but a vector dimension of $T_{x}T_{y}\gg T_{y}$. This has the benefit of limiting the effect of $g$ in BlockLDLQ’s error bound $gm\mu^{2}\sigma^{2}\mbox{tr}(H^{1/2})^{2}/n$ while achieving a high dimension for TCQ. Algorithm 5 in the Appendix describes this in more detail.
4.1 Lookup-Free Computed Codes
Table 3: Wikitext2 and C4 perplexity ($\downarrow$), ctx. 4096, QTIP with pure-computed codes. Even without fine-tuning, pure-computed QTIP outperforms QuIP##\## with fine-tuning at almost all models and sizes.
4 Bit No FT $\approx$4 Bit 3 Bit No FT $\approx$3 Bit 2 Bit No FT $\approx$2 Bit
FP16 1MAD 3INST QuIP# QuIP# AQLM 1MAD 3INST QuIP# QuIP# AQLM 1MAD 3INST QuIP# QuIP# AQLM
W2 5.12 5.17 5.17 5.22 5.19 5.21 5.38 5.40 5.60 5.41 5.46 7.05 6.82 8.22 6.19 6.64
2-7 C4 6.63 6.71 6.71 6.79 6.75 6.75 6.99 7.01 7.34 7.04 7.08 9.14 8.96 11.0 8.16 8.56
W2 4.57 4.62 4.62 4.65 4.63 4.64 4.75 4.74 4.90 4.78 4.83 5.59 5.52 6.06 5.35 5.70
2-13 C4 6.05 6.10 6.10 6.15 6.13 6.14 6.28 6.28 6.50 6.35 6.37 7.46 7.39 8.07 7.20 7.59
W2 3.12 3.16 3.16 3.18 3.18 3.19 3.27 3.27 3.41 3.35 3.36 3.87 3.90 4.16 3.91 3.94
2-70 C4 4.97 5.00 5.00 5.02 5.02 5.03 5.09 5.09 5.20 5.15 5.17 5.70 5.69 6.01 5.71 5.72
Here, we use 1MAD and 3INST with $L=16,V=1,T_{x}=T_{y}=16$. Setting $T_{x}=T_{y}=16$ enables using a $16\times 16$ MMA tile per trellis sequence to perform matrix multiplication during inference. $16\times 16$ MMA tiles form the basis of many types of “AI hardware,” making fast decoding relatively simple [7]. We do not perform fine-tuning since the codes themselves are not tunable, but these codes are fully compatible with QuIP##\##-style fine-tuning (recall that QuIP##\##’s codebook is also not tunable). Table 3 shows that both 1MAD and 3INST significantly outperform QuIP##\## without fine-tuning (AQLM does not have numbers without fine-tuning). Even at 4 bits, where all methods are close to lossless, QTIP results in significant improvements. Notably, the computed-code QTIP variants without fine-tuning outperforms both QuIP##\## and AQLM with fine-tuning on almost all models and sizes, showing that fine-tuning is not a silver bullet.
4.2 Hybrid Lookup-Computed Codes
Table 4: Wikitext2 and C4 perplexity ($\downarrow$), QTIP with the hybrid-computed code. QTIP enables high-dimensional quantization and outperforms state-of-the-art vector quantization approaches.
ctx. 2048, X = GPTVQ, Y = 0.13 ctx. 4096, X = AQLM, Y $\approx$ 0
Wiktext2 C4 Wikitext2 C4
Method Bits 1-7 1-13 1-30 1-65 1-7 1-13 1-30 1-65 2-7 2-13 2-70 2-7 2-13 2-70
FP16 16.0 5.68 5.09 4.10 3.53 7.04 6.61 5.98 5.62 5.12 4.57 3.12 6.63 6.05 4.97
X 4+Y 5.94 5.20 4.18 3.64 5.21 4.65 3.19 6.75 6.14 5.03
QuIP# 4.00 5.76 5.17 4.18 3.60 7.18 6.67 6.03 5.66 5.19 4.63 3.18 6.75 6.13 5.02
QTIP 4.00 5.72 5.15 4.15 3.58 7.13 6.65 6.01 5.64 5.17 4.61 3.16 6.69 6.09 5.00
X 3+Y 6.32 5.31 4.38 3.79 5.46 4.82 3.36 7.08 6.37 5.17
QuIP# 3.00 5.98 5.31 4.36 3.70 7.39 6.83 6.17 5.77 5.41 4.78 3.35 7.04 6.35 5.15
QTIP 3.00 5.85 5.24 4.26 3.68 7.26 6.74 6.09 5.71 5.29 4.71 3.26 6.88 6.23 5.08
X 2+Y 9.64 6.58 5.63 4.91 6.64 5.65 3.94 8.56 7.51 5.72
QuIP# 2.00 6.86 5.97 5.02 4.36 8.36 7.48 6.71 6.19 6.19 5.35 3.91 8.16 7.20 5.71
QTIP 2.00 6.52 5.80 4.83 4.21 7.99 7.31 6.56 6.08 5.91 5.26 3.78 7.76 6.99 5.56
Table 5: Zeroshot accuracy ($\uparrow$), QTIP with the hybrid-computed code.
2-70 2-13 2-7
Mthd. Bits ArcC ArcE PiQA Wino Bits ArcC ArcE PiQA Wino Bits ArcC ArcE PiQA Wino
FP16 16 51.1 77.7 81.1 77 16 45.6 73.3 73.5 69.6 16 40 69.3 78.5 67.3
AQLM 4.07 51.0 78.1 81.4 76.9 3.94 43.9 72.2 78.6 70.4 4.04 40.3 68.9 77.7 67.3
QuIP# 4 50.6 78.1 81.4 77.1 4.00 45.5 73.9 78.9 69.9 4.00 40.5 69.1 78.4 67.6
QTIP 4 50.0 77.6 81.5 77.0 4.00 45.9 73.2 78.6 69.9 4.00 40.4 69.2 78.5 67.4
AQLM 3.01 50.0 77.6 81.3 77.2 3.03 43.6 73.5 77.8 67.6 3.04 38.7 67.8 76.6 68.4
QuIP# 3 50.0 77.7 81.4 76.4 3.00 44.0 72.5 78.4 69.1 3.00 39.2 68.4 77.3 66.5
QTIP 3 49.8 77.5 81.2 76.3 3.00 43.6 72.5 78.2 69.5 3.00 39.8 69.7 78.0 66.8
AQLM 2.07 47.9 77.7 80.4 75.9 1.97 38.5 67.0 75.1 69.5 2.02 33.6 62.8 73.5 64.6
QuIP# 2 48.7 77.3 80.3 75.9 2.00 39.5 69.3 77.3 67.7 2.00 34.6 64.6 75.1 64.9
QTIP 2 48.1 76.9 80.1 76.5 2.00 39.2 70.6 77.8 71.0 2.00 35.3 63.9 75.3 66.7
Table 6: QTIP vs. QuIP##\##, Llama 3 (ctx. 8192 for perplexity). Although Hessian-based rounding generally underperforms on Llama 3, the focus of this work is on what to quantize with (TCQ vs. VQ). Here, the high-dimensionality of TCQ in QTIP improves over the low-dimensional VQ in QuIP##\##.
3-70 ppl ($\downarrow$) 3-70 zeroshot acc ($\uparrow$) 3-8 ppl ($\downarrow$) 3-8 zeroshot acc ($\uparrow$)
Mthd. Bits W2 C4 ArcC ArcE BoolQ PiQA Wino W2 C4 ArcC ArcE BoolQ PiQA Wino
FP16 16.0 2.59 5.78 60.5 86.9 85.3 82.4 80.3 5.54 7.10 50.2 80.1 81.0 79.7 72.9
QuIP# 4.00 2.99 5.96 35.0 67.3 84.7 71.9 76.7 5.81 7.32 50.2 79.7 81.3 79.7 73.1
QTIP 4.00 2.75 5.83 56.1 83.9 85.8 81.3 80.6 5.67 7.20 50.2 79.6 79.5 79.4 73.4
QuIP# 3.00 3.59 6.18 31.1 36.6 85.7 58.8 76.4 6.27 7.71 46.4 77.4 79.9 77.9 72.9
QTIP 3.00 3.18 5.98 48.6 77.8 85.0 77.8 79.7 6.01 7.48 49.2 79.3 80.0 79.2 74.5
QuIP# 2.00 5.77 7.46 18.3 32.2 82.1 54.7 68.9 7.84 9.06 39.2 72.9 76.6 75.6 68.2
QTIP 2.00 4.97 6.80 28.0 35.2 83.6 57.1 72.6 7.33 8.62 44.2 75.2 76.7 77.6 70.7
Here, we use the hybrid lookup-computed code with $L=16,V=2,T_{x}=T_{y}=16,Q=9$. Setting $Q=9$ gives a 2KiB codebook, which fits in L1 cache even after duplication for bank conflicts (32$\times$) on modern GPUs. This codebook is differentiable, so we can fine-tune it: to evaluate this, we fine-tune using QuIP##\##’s methodology, tuning both the codebook entries and the as-yet-unquantized weights in a blockwise fashion. Table 4 shows the perplexity of quantized Llama 1 and 2 models. In all cases, QTIP outperforms the other vector quantization-based methods. Even at 3 and 4 bits, where QuIP##\## and AQLM are close to lossless, QTIP roughly halves the perplexity gap. These results also show the importance of dimensionality. Note that the 3- and 4-bit Llama 2 70B numbers here match those in 3. Since Table 3 uses a pure-computed code without fine-tuning, fine-tuning has no effect in these regimes and the improvement over QuIP##\## is purely from dimensionality.
Table 5 shows zeroshot results computed with LM Eval, which are slightly random; QTIP generally matches or exceeds QuIP##\## and AQLM on these tasks [16]. Table 6 contains results on Llama 3. Like other works, we have observed that Hessian-based adaptive rounding (GPTQ/LDLQ/BlockLDLQ) performs worse than expected on Llama 3 [17]. Since the contribution and focus of this work is what to round with (TCQ) and not how to round (BlockLDLQ), we only compare against the proximal baseline QuIP##\##, which uses BlockLDLQ with VQ. QTIP significantly improves upon QuIP##\## at all model sizes and bitrates, once again showing the dimensionality advantage of TCQ over VQ.
4.3 Inference Speed
Table 7: RTX4090 (1TB/s) Throughput
2-7B 2-70B
Mthd. tok/s
matvec
mem bw
tok/s
matvec
mem bw
FP16 33.1 OOM OOM
AQLM 20.6 8.27
QuIP# 106.3 697 GB/s 25.9 867 GB/s
QTIP 105.2 628 GB/s 25.8 840 GB/s
Table 7 shows the batch size 1 inference speed of QTIP, QuIP##\##, and AQLM on Llama 2 7b and 70b without matrix fusion. Here, the design choices of QTIP and QuIP##\## become apparent. Whereas AQLM uses a codebook that is too large to fit in cache and thus prevents fast inference, both QTIP and QuIP##\## achieve significant speedups over FP16. According to NVIDIA Nsight Systems, the matrix-vector multiply kernels of QuIP##\## and QTIP both run at near maximum memory bandwidth [29]. This means that a significant part of the gap between their kernel memory bandwidths and effective tok/s is due to poor kernel fusion in the rest of the inference pipeline, which could easily be optimized to further improve throughput. Finally, while it is impressive that both QuIP##\## and QTIP are 3-5$\times$ faster than AQLM, it is even more impressive that QTIP is able to match QuIP##\##’s throughput with an effective dimension size of 256, or 32×32$\times$32 × larger than QuIP##\##’s. This means that the improved quantization quality of QTIP comes with no additional inference-time cost.
5 Conclusion
We present QTIP, a weight-only post-training quantization algorithm that achieves state-of-the-art results through the use of trellis-coded quantization (TCQ). TCQ enables tractable ultra-high dimensional quantization, significantly reducing quantization distortion over vector quantization (VQ). However, naive TCQ does not admit fast inference due to sequential bottlenecks during decoding and needing to store a large codebook. QTIP solves this problem through a novel combination of incoherence processing, the hardware-efficient bitshift trellis, and fast computed codes. Specifically, QTIP introduces a series of compute-based pseudorandom Gaussian codes that, when used in conjunction with the bitshift trellis and incoherence processing, simultaneously achieves state-of-the-art PTQ quality and fast inference. QTIP improves quantization quality at all tested bitrates over the latest VQ-based PTQ methods, QuIP##\## and AQLM, further pushing the boundary of LLM PTQ. QTIP’s codes use as few as 2 instructions per weight during decoding, enabling matrix-vector multiplication to run at over 80% of peak memory bandwidth on modern GPUs. Altogether, our results indicate that high dimensional quantization is necessary for high-quality compression, and QTIP is the first LLM PTQ method to scale to ultra-high dimensions while supporting fast inference.
Appendix A Appendix
A.1 Additional Results
Table 8: Zeroshot results for the 1MAD code.
Bits ArcC (acc) ArcE (acc) BoolQ (acc) PiQA (acc) Wino (acc)
2-7 16 39.9 69.3 71.1 78.4 67.2
2-7 4 39.0 69.4 72.0 78.4 67.9
2-7 3 38.8 68.0 68.2 77.6 68.4
2-7 2 32.1 63.5 66.3 73.3 62.7
2-13 16 45.6 73.3 69.1 78.7 69.7
2-13 4 45.6 72.9 68.1 78.7 70.3
2-13 3 42.2 71.0 69.9 78.6 69.8
2-13 2 38.5 71.5 71.4 75.9 68.9
2-70 16 51.2 77.7 76.7 81.1 76.9
2-70 4 51.1 77.8 75.2 81.5 77.0
2-70 3 50.8 77.8 77.9 80.7 76.3
2-70 2 49.3 77.7 83.3 80.4 75.7
Table 9: Zeroshot results for the 3INST code.
Bits ArcC (acc) ArcE (acc) BoolQ (acc) PiQA (acc) Wino (acc)
2-7 16 39.9 69.3 71.1 78.4 67.2
2-7 4 40.2 68.5 70.3 78.0 67.7
2-7 3 40.2 68.6 73.0 77.5 65.4
2-7 2 32.9 61.9 65.5 74.5 65.0
2-13 16 45.6 73.3 69.1 78.7 69.7
2-13 4 45.4 72.7 67.9 78.5 69.9
2-13 3 44.5 72.6 70.1 78.5 69.4
2-13 2 38.7 68.2 63.6 75.6 68.7
2-70 16 51.2 77.7 76.7 81.1 76.9
2-70 4 50.3 77.9 77.3 80.7 76.5
2-70 3 50.9 78.3 78.8 81.1 77.5
2-70 2 48.0 76.5 76.7 80.1 77.6
Table 10: Llama 1 Zeroshot results for the Hybrid code
Bits ArcC (acc) ArcE (acc) BoolQ (acc) PiQA (acc) Wino (acc)
1-7 16 38.2 67.4 73.1 78.4 67.0
1-7 4 38.8 67.1 74.2 78.3 67.1
1-7 3 37.0 65.7 74.1 77.7 67.3
1-7 2 35.3 64.9 72.9 76.1 65.4
1-13 16 43.9 74.6 68.5 78.8 70.1
1-13 4 43.4 73.7 68.2 79.1 70.1
1-13 3 42.2 74.2 68.0 78.7 70.5
1-13 2 39.7 72.1 66.6 77.6 68.9
1-30 16 46.7 75.4 68.4 81.0 72.6
1-30 4 46.7 75.4 69.9 81.0 73.3
1-30 3 47.8 75.0 70.0 80.4 73.6
1-30 2 44.0 72.7 72.8 78.7 71.7
1-65 16 47.0 75.3 82.3 81.5 77.2
1-65 4 46.8 74.5 82.8 81.4 76.6
1-65 3 46.8 75.3 83.0 81.3 75.9
1-65 2 44.4 74.2 83.1 80.4 75.7
A.1.1 Lookup-Only Codes
Table 11: Wikitext2 and C4 perplexity ($\downarrow$), ctx. 4096, QTIP with a size $2^{14}$ LUT codebook. This codebook is too large (32KB) for current GPU L1 caches, but could fit on near-future hardware.
$\sim$4 Bit $\sim$3 Bit $\sim$2 Bit
FP16 QTIP QuIP# AQLM QTIP QuIP# AQLM QTIP QuIP# AQLM
W2 5.12 5.16 5.19 5.21 5.30 5.41 5.46 5.89 6.19 6.64
2-7 C4 6.63 6.68 6.75 6.75 6.86 7.04 7.08 7.78 8.16 8.56
W2 3.12 3.15 3.18 3.19 3.26 3.35 3.36 3.77 3.91 3.94
2-70 C4 4.97 4.99 5.02 5.03 5.07 5.15 5.17 5.55 5.71 5.72
Table 12: Wikitext2 and C4 zeroshot accuracy ($\uparrow$), QTIP with a size $2^{14}$ LUT codebook. This codebook is too large (32KB) for current GPU L1 caches, but could fit on near-future hardware.
Bits ArcC (acc) ArcE (acc) BoolQ (acc) PiQA (acc) Wino (acc)
2-7 16.0 40.0 69.3 71.0 78.5 67.3
2-7 4.00 40.3 69.2 73.0 78.1 67.5
2-7 3.00 39.1 69.3 69.6 77.8 66.3
2-7 2.00 37.0 64.6 67.2 75.6 66.9
2-70 16.0 51.1 77.7 76.6 81.1 77.0
2-70 4.00 50.1 77.5 76.4 81.3 77.3
2-70 3.00 50.6 77.9 78.0 81.1 76.1
2-70 2.00 47.1 76.9 79.5 80.1 76.3
Here, we use a pure-lookup code $\sim\mathcal{N}(0,1)$ with $L=14,V=1,T_{x}=32$, $T_{y}=8$, and QuIP##\##’s fine-tuning scheme. These parameters show what performance QTIP could achieve if we did not care about fast inference today. Specifically, a pure-lookup codebook is tunable, and setting $T_{y}=8$ reduces the BlockLDLQ group size while maintaining high dimensionality (256). This codebook uses 32KB; this only fits in GPU L1 cache with bank conflicts. Setting $T_{x}=32,T_{y}=8$ corresponds to using a larger MMA tile size than current GPUs allow for. The largest tile size is usually 16 in the $T_{x}$ dimension, meaning that a $32\times 8$ trellis needs two tiles. Thankfully, hardware required to serve such a model quickly is likely only a few years away, as these parameters are only slightly outside of what today’s hardware is capable of.
Table 11 shows that QTIP outperforms both QuIP##\## and AQLM at all compression ratios, with 3 bit QTIP achieving similar quality as 4 bit AQLM. While it is not fair to compare this QTIP setup with QuIP##\##, since QuIP##\## was designed for fast inference, we note that AQLM’s VQ codebook uses $12^{16}\times 8\times 2=$ 1 MiB. This is 32 times larger than the QTIP codebook here, and would require 32 MiB of L1 cache to read from without bank conflicts. Not only is this orders of magnitude larger than current L1 caches (256KB on the H100), it is even larger than many L2 caches!
A.2 QTIP with BlockLDLQ
Here, we detail how we use TCQ within BlockLDLQ to produce our experimental setup. Essentially, QTIP is used as a high dimensional $T_{x}T_{y}$ quantizer within BlockLDLQ and is a drop-in replacement for vector quantization in BlockLDLQ. The regular blockLDLQ step $Q(W+(W-\hat{W})A)$ is exactly the same, and the only difference is in how $Q$ rounds. Instead of rounding each row of $x=W+(W-\hat{W})A$ independently, it groups $T_{x}$ rows into a block to round as $m/T_{x}$ high-dimensional sequences.
Algorithm 5 QTIP with BlockLDLQ
0:  $W\in\mathbb{R}^{m\times n},H\in\mathbb{R}^{n\times n},T_{x},T_{y},L,k,V,$ code $C$.
  $\hat{W}\leftarrow 0_{m,n}$
  $LDL^{T}\leftarrow T_{y}$-block LDL decomposition of $H$
  $A\leftarrow L-I$
  for $j\in\{n/T_{y}-1,n/T_{y}-2,...,0\}$ do
     $x\leftarrow W_{:,jT_{y}:(j+1)T_{y}}+(W_{:,jT_{y}:}-\hat{W}_{:,jT_{y}:})A_{jT_{y}:,jT_{y}:(j+1)T_{y}}$
     $x\leftarrow x\text{.reshape}(m/T_{x},T_{x}T_{y})$
     $\hat{x}\leftarrow\text{Viterbi}(x,(L,k,V)\text{ bitshift trellis},C)$ (row-wise)
     $\hat{W}_{:,jT_{y}:(j+1)T_{y}}\leftarrow\hat{x}\text{.reshape}(m,T_{y})$
  end for
  Quantized $\hat{W}$.
A.3 Implementation Details
A.3.1 Code
Our code is available at https://github.com/Cornell-RelaxML/qtip.
A.3.2 Hessian Generation
Hessian matrices were generated with 6144 sequences of length 2048 for Llama 1, 4096 for Llama 2, and 4096 sequences of 6192 for Llama 3. All sequences were sampled from the RedPajama dataset [8].
A.3.3 Hadamard Matrices
We use Hadamard matrices from Neil Sloane’s website [32].
A.3.4 Perplexity
We use the OPTQ “Wikitext2” and “C4” test splits to calculate perplexity [15].
A.3.5 Fine Tuning
We use the same fine-tuning hyperparemters as QuIP##\##, except that we do not perform matrix fusion [35]. This results in having to quantize 7 matrices (q, k, v, o, up, gate, down) instead of 4 (qkv, o, upgate, down). We quantize and fine tune in the following order: (v, q, k, o, up gate, down).
A.4 Example Generations
The following are some sample generations from Llama 2 70B (not instruction tuned) quantized to 2 bits with QTIP, truncated to 256 tokens.
Prompt: It is a truth universally acknowledged QTIP: It is a truth universally acknowledged that a single man in possession of a good fortune must be in want of a wife. However little known the feelings or views of such a man may be on his first entering a neighbourhood, this truth is so well fixed in the minds of the surrounding residents, that he is considered the rightful property of some one or other of their daughters. “My dear Mr. Bennet,” said his lady to him one day, “have you heard that Netherfield Park is let at last?” “But it is,” returned she; “for Mrs. Long has just been here, and she told me that Netherfield is taken by a young man of large fortune from the north of England; that he came down on Monday in a chaise and four to see the place, and was so much delighted with it that he agreed with Mr. Morris immediately; that he is to take possession before Michaelmas, and that he is now in town in search of a wife.” “I see no occasion for that. You and the girls may go, or you may send them by themselves, which do you prefer? Only let me know what you are meaning to do, for I shall certainly not be at home. You
Prompt: Call me Ishamel QTIP: Call me Ishmael. Some years ago—never mind how long precisely—having little or no money in my purse, and nothing particular to interest me on shore, I thought I would sail about a little and see the watery part of the world. It is a way I have of driving off the spleen, and regulating the circulation. Whenever I find myself growing grim about the mouth; whenever it is a damp, drizzly November in my soul; whenever I find myself involuntarily pausing before coffin warehouses, and bringing up the rear of every funeral I meet; and especially whenever my hypochondriacal palate is out of sorts, and I am desperately dull, and in the dumps, and have such an overpowering sense of fatality that I find myself involuntarily gravitating toward Cemeteries—and especially toward those open ones which are occasionally cheered by the gleam of a bonfire, or a cigar, or a pipe, or a lantern, or a candle, or a torch, or the moon, I then take to the ship. There is nothing surprising in this,